Use the definition of limit to prove that the sequence {n −(1/n)}n=1 to infinity is divergent.


1
Expert's answer
2021-02-01T05:40:21-0500

we need to show that given M "\gt 0," there exists a natural number N(depending on M) such that "(n-\frac{1}{n})n \gt M" whenever "n \gt N"

now,

"(n-\frac{1}{n})n \gt n-1 \gt M" whenever "n \gt N = M+1"

therefore, taking N=M,

"(n-\frac{1}{n})n \gt M" whenever n "\gt N"